# Walk Before You Dance: High-fidelity and Editable Dance Synthesis via Generative Masked Motion Prior

### v2

@@ 3 Proposed Method: DanceMosaic
Given an input music signal AA, a text prompt $T$, and pose constraints on specific joints or body parts $P$, our goal is to synthesize a physically plausible 3D dance sequence that aligns with the music and text prompts, while satisfying the pose constraints. To achieve this, we propose a multi-tower masked motion model, which incorporates music and pose guidance modules into the text-conditioned masked motion model. The overview of DanceMosaic, shown in Fig. 2, consists of a dance motion tokenizer and three integrated masked transformers, including a Text-to-Motion (T2M) model, Music Tower (MT), and Pose Tower (PT). Motion tokenizer transforms the raw motion sequence into categorical discrete motion tokens within a learned codebook (Sec 3.1). The T2M model learns to predict masked motion tokens, conditioned on textual prompts, like dance genres and action descriptions. The music and pose towers in Sec. 3.2 are designed to introduce the influence of music and pose signals to the T2M model via synchronized and progressive multimodal training (Sec. 3.3). During inference, the multimodal guidance modules, including classifier-free logits guidance and inference-time token optimization (ITTO), steer the token sampling process so that generated dance sequence is aligned with diverse prompt signals (Sec. 3.4).
@@ 3.2 Multi-tower Masked Motion Model
Guidance Signals and Per-frame Alignment. For the music tower, we adopt a frozen Jukebox model (dhariwal2020jukebox) to provide a per-frame music embedding sequence, enabling precise synchronization between dance movements and musical rhythm. For pose tower, we introduce a spatial editing signal ${P}\in\mathbb{R}^{N\times J\times 3}$ into the masked token reconstruction, allowing targeted modifications to upper-body, lower-body or joint-specific movements. The pose signal P{P} encodes the $(x,y,z)$ positions of each of the $J$ skeleton joints over NN frames. A subset of joint locations is provided as spatial constraints and remains fixed during generation, while the positions of the remaining joints, which are set to zero, are editable. To effectively integrate these guidance signals, we leverage learnable music and pose encoders to transform the respective inputs into latent tokens matching the dimensionality of motion tokens. These music and pose latent tokens are then precisely aligned with motion tokens on a per-frame basis via token-wise addition, ensuring coherent synchronization across dance, music, and pose inputs.
@@ 3.3 Synchronized Progressive Masked Training
Synchronized Masking. The music and pose-guidance towers are trained using synchronized masked training, where both the T2M model and guidance towers receive identically masked motion token. This synchronized masking ensures that the guidance towers can effectively leverage the motion distributions learned by the T2M model. The music tower is trained to recover the masked motion tokens, conditioned on the text prompt $T$, the music signal AA, and the corrupted motion token sequence $Z_{\mathbf{M}}$ by minimizing the CE loss:
@@
Progressive Training. Integrating music, text, and pose control signals into a dance generation model is challenging due to conflicting loss functions that create competing gradients during training. In particular, the kinematic terms in the music tower loss ($\mathcal{L}_{MT}$) aim to recover the full body skeleton joints, while the discrepancy term in the pose tower loss ($\mathcal{L}_{PT}$) focuses on restoring only the non-editable joints. To address this, we adopt multimodal progressive training, which incrementally incorporates modality-specific signals. This approach allows each modality of the model to be trained according to its respective loss function, mitigating gradient conflicts. The training process begins with training T2M model on text-to-motion dataset using textual prompts that describe a broad range of human actions. Next, we introduce the music tower, trained on dance-only datasets while keeping the T2M model frozen. This stage incorporates both music and genre prompts. At last, pose tower is trained with the frozen T2M using pose and genre prompts. Although the pose tower is not explicitly trained with music signal AA, integrating the music and pose towers with the T2M allows the pose $P$ and music AA signals to simultaneously manipulate the prior motion distribution learned by the T2M.
@@ 4 Experiments
Methods Top-1 $\uparrow$ FID $\downarrow$ $M$ Dist $\downarrow$ Diversity $\uparrow$

### v3

@@ 3 Proposed Method: DanceMosaic
Given an input music signal $A$, a text prompt $T$, and pose constraints on specific joints or body parts $P$, our goal is to synthesize a physically plausible 3D dance sequence that aligns with the music and text prompts, while satisfying the pose constraints. To achieve this, we propose a multi-tower masked motion model, which incorporates music and pose guidance modules into the text-conditioned masked motion model. The overview of DanceMosaic, shown in Fig. 2, consists of a dance motion tokenizer and three integrated masked transformers, including a Text-to-Motion (T2M) model, Music Tower (MT), and Pose Tower (PT). Motion tokenizer transforms the raw motion sequence into categorical discrete motion tokens within a learned codebook (Sec 3.1). The T2M model learns to predict masked motion tokens, conditioned on textual prompts, like dance genres and action descriptions. The music and pose towers in Sec. 3.2 are designed to introduce the influence of music and pose signals to the T2M model via synchronized and progressive multimodal training (Sec. 3.3). During inference, the multimodal guidance modules, including classifier-free logits guidance and inference-time token optimization (ITTO), steer the token sampling process so that generated dance sequence is aligned with diverse prompt signals (Sec. 3.4).
@@ 3.2 Multi-tower Masked Motion Model
Guidance Signals and Per-frame Alignment. For the music tower, we adopt a frozen Jukebox model (dhariwal2020jukebox) to provide a per-frame music embedding sequence, enabling precise synchronization between dance movements and musical rhythm. For pose tower, we introduce a spatial editing signal ${P}\in\mathbb{R}^{N\times J\times 3}$ into the masked token reconstruction, allowing targeted modifications to upper-body, lower-body or joint-specific movements. The pose signal P{P} encodes the $(x,y,z)$ positions of each of the $J$ skeleton joints over $N$ frames. A subset of joint locations is provided as spatial constraints and remains fixed during generation, while the positions of the remaining joints, which are set to zero, are editable. To effectively integrate these guidance signals, we leverage learnable music and pose encoders to transform the respective inputs into latent tokens matching the dimensionality of motion tokens. These music and pose latent tokens are then precisely aligned with motion tokens on a per-frame basis via token-wise addition, ensuring coherent synchronization across dance, music, and pose inputs.
@@ 3.3 Synchronized Progressive Masked Training
Synchronized Masking. The music and pose-guidance towers are trained using synchronized masked training, where both the T2M model and guidance towers receive identically masked motion token. This synchronized masking ensures that the guidance towers can effectively leverage the motion distributions learned by the T2M model. The music tower is trained to recover the masked motion tokens, conditioned on the text prompt $T$, the music signal $A$, and the corrupted motion token sequence $Z_{\mathbf{M}}$ by minimizing the CE loss:
@@
Progressive Training. Integrating music, text, and pose control signals into a dance generation model is challenging due to conflicting loss functions that create competing gradients during training. In particular, the kinematic terms in the music tower loss ($\mathcal{L}_{MT}$) aim to recover the full body skeleton joints, while the discrepancy term in the pose tower loss ($\mathcal{L}_{PT}$) focuses on restoring only the non-editable joints. To address this, we adopt multimodal progressive training, which incrementally incorporates modality-specific signals. This approach allows each modality of the model to be trained according to its respective loss function, mitigating gradient conflicts. The training process begins with training T2M model on text-to-motion dataset using textual prompts that describe a broad range of human actions. Next, we introduce the music tower, trained on dance-only datasets while keeping the T2M model frozen. This stage incorporates both music and genre prompts. At last, pose tower is trained with the frozen T2M using pose and genre prompts. Although the pose tower is not explicitly trained with music signal $A$, integrating the music and pose towers with the T2M allows the pose $P$ and music $A$ signals to simultaneously manipulate the prior motion distribution learned by the T2M.
@@ 4 Experiments
Methods Top-1 $\uparrow$ FID $\downarrow$ MM Dist $\downarrow$ Diversity $\uparrow$